# Language-guided Open-world Video Anomaly Detection

### v1

@@ 6.2 Comparison with State-of-the-Arts
As demonstrated in Figs. 2, 2 and 3, our method achieves state-of-the-art performance in both zero-shot detection and classification. In Tab. 2, LaGoVAD outperforms existing approaches, including the LLM-based method [40], vision-language aligned methods [35, 21], and the multi-domain generalization method [9] on detection. Notably, on XD-Violence, our approach achieves a 20% improvement over prior methods. For fair comparison with OVVAD [34], which leverages AIGC-augmented training, we construct a comparable set and still observe superior performance. In Tab. 3, empirical results demonstrate that our approach outperforms both the CLIP baseline and state-of-the-art action recognition models under identical CLIP variants, attributable to its capacity for selectively attending to abnormal patterns in videos. While sharing the same feature extractor and a similar two-branch architecture capable of detection and classification, our framework achieves comprehensive improvements over VadCLIP [35]. Notably, it exhibits a 27% gain in detection performance (measured by AP) and a 68% improvement in classification accuracy on XD-Violence.
Table 4: Ablation on each component. guided refers to guiding the detection with language. Det. Avg. refers to the average zero-shot detection performance on seven datasets. Cls. Avg. refers to the average zero-shot classification performance on UCF-Crime and XD-Violence.
$\mathcal{L}_{\text{dvs}}$ $\mathcal{L}_{\text{neg}}$ guided PreVAD Det. Avg. Cls. Avg.
@@
6.3 Ablation Studies
Dataset Effectiveness
To quantify dataset impacts, we compare VadCLIP [35] trained by three datasets in Tabs. 2 and 3 via average metrics in detection and classification. Experimental results reveal that the model trained with PreVAD outperforms the one trained with UCF-Crime by 14% in detection (average metrics on six other datasets) and 88% in classification (average metrics on XD-Violence) while surpassing the one trained with XD-Violence by 7.6% in detection and 44% in classification, respectively. This substantial margin validates that a larger and more diverse dataset can significantly improve zero-shot performance.
Architecture Effectiveness
As demonstrated in Tabs. 2 and 3, when trained on identical datasets, our LaGoVAD framework achieves consistent improvements over VadCLIP, with gains of 7.2% in average detection performance on seven datasets and 2.8% in classification across two datasets. This confirms the superiority of our paradigm in open-world scenarios.
Module Effectiveness
We conduct a series of ablations to validate each module of our model in Tab. 4. Removing either the dynamic video synthesis module or the contrastive learning with hard negative mining led to a noticeable degradation in detection and classification performance. When both are removed, the model exhibits a significant decline in zero-shot performance. We also validate the effect of language guidance by removing it in our experiments. When disabling the language guidance, we followed approaches in [35, 34] to place the fusion module after the detection stage, which does not condition detection results on the given text. Experiment shows that without language guidance, in-domain performance only slightly dropped, while cross-domain performance decreased significantly. It indicates that the conventional paradigm lack the capacity to incorporate user-defined guidance for detection, thereby limiting their adaptability to open-world scenarios.
@@ 6.4 Qualitative Results
Figure 7 visualizes the prediction results of PEL [21], VadCLIP [35] and our proposed LaGoVAD under concept drift. As illustrated in Fig. 7a, the video depicts an incident of stealing a cattle, which is considered as abnormal in the UCF-Crime dataset. When users exclude such categories (e.g., in traffic or violence detection scenarios), our method could dynamically adapt to it by redefining anomalies. In Fig. 7b, the video shows a dog knocking over a trashcan, labeled as normal in UCF-Crime. However, when users explicitly define this behavior as an anomaly, our method seamlessly adjusts its detection logic. In contrast, the two other methods are not able to address concept drift under different user requirements. These results validate the practicability and effectiveness of our paradigm in open-world video anomaly detection. More ablations and visualizations are in the supp (Sec. F).
Figure 7: Visualization of predicted anomaly scores for two cases where concept drift occurs in open-world scenarios, compared with VadCLIP [35] and PEL [21] trained on UCF-Crime [25]. Both videos are from the UCF-Crime dataset, where (a) is labeled as abnormal and (b) is labeled as normal.

### v2

@@ 6.2 Comparison with State-of-the-Arts
Under the comprehensive evaluation of Protocol 1 (Tabs. 2,4), our approach surpasses others across all datasets, which includes comparisons with related methods in open-vocabulary setting (OVVAD, ActionCLIP, ViFi-CLIP), cross-domain setting (MultiDomain), or scene-dependent setting (CMRL). Notably, on XD-Violence, it achieves improvements of 20% and 32% in detection and classification, respectively. Under the concept drift evaluation of Protocol 2 (Tab. 4), LaGoVAD achieves better performance than multi-modal methods and LLM-based methods, while avoiding the significant computational overhead from huge parameters.
Figure 4: Visualization of different methods under concept drift. Knocking over a trashcan is considered normal in (a) but abnormal in (b). All models are prompted with the corresponding definition.
6.3 Ablation Studies
Dataset & Architecture Effectiveness. To quantify dataset and architecture impacts, we compare VadCLIP trained on PreVAD. The results reveal that the model trained on PreVAD outperforms the one trained on UCF-Crime by 14% in detection (average metric on six other datasets) and 88% in classification (average metric on XD-Violence) while also surpassing the one trained on XD-Violence by 7.6% in detection and 44% in classification, respectively. This substantial margin validates that a larger and more diverse dataset can significantly improve zero-shot performance. When trained on PreVAD, our LaGoVAD achieves consistent improvements over VadCLIP, with gains of 7.2% in average detection performance across seven datasets and 2.8% in classification across on two datasets. This confirms the superiority of our approach in open-world scenarios.
Table 5: Ablation on each component. Lang-guided: language guiding. Det. Avg.: average zero-shot temporal detection performance on seven datasets. Cls. Avg.: average zero-shot multi-classification performance on UCF and XD.
Lang-guided $\mathcal{L}_{\text{dvs}}$ $\mathcal{L}_{\text{neg}}$ PreVAD Det. Avg. Cls. Avg.
@@ 6.3 Ablation Studies
Module Effectiveness. As shown in Tab. 5, removing either $\mathcal{L}_{\text{dvs}}$ or $\mathcal{L}_{\text{neg}}$ led to a noticeable degradation in detection and classification performance. When both are removed, the model exhibits a significant decline in zero-shot performance. When disabling the language guidance, we followed approaches in (Wu et al., 2024c; b) to place the fusion module after the detection stage, which does not condition detection results on the given text. Experiment shows that without language guidance, cross-domain performance decreased significantly, which indicates that the conventional paradigm lack the capacity to incorporate user-defined guidance for detection, thereby limiting their adaptability to open-world scenarios. More ablations of modules and data are in the supp Sec. F.
@@ 6.4 Qualitative Results
Fig. 4 visualizes the performance under concept drift. The conventional method (VadCLIP) fails to handle dynamic definitions, producing same scores under the training definition. Although LLM based methods (LAVAD, Qwen2.5-VL) can take definition prompts, LAVAD fails to recognize the anomaly due to its limited understanding of dynamic events. Qwen2.5-VL recognizes it but cannot localize it precisely. Our method, in contrast, adapts to the dynamic definition and achieves precise anomaly localization.
@@ 7 Conclusion
Appendix A Limitation and Future Work
While our work introduces a novel paradigm for addressing concept drift in open-world video anomaly detection, we acknowledge that limitation exists. The current architecture instantiates the paradigm through a simplified design, leaving room for architectural refinements to better capture temporal dependencies and multimodal interactions. For future work, researchers may establish comprehensive benchmarks beyond the current evaluation to systematically evaluate prompt compliance and open-world capability. Moreover, through our proposed pipeline, a larger dataset could be collected to further boost the performance.
Appendix B Details of LaGoVAD
B.1 Architecture
Dynamic Video Synthesis
In practical scenarios, abnormal events typically occupy a small proportion of the total video duration. However, since people often edit videos to highlight events of interest before uploading them to the Internet, web-sourced videos generally exhibit high anomaly ratios. For instance, the test sets of MSAD, PreVAD, and LAD contain 42%, 39%, and 38% of videos, respectively, where abnormal events occupy over 70% of the total duration. This data characteristic leads to the loss of normal contextual information, consequently hindering the model’s ability to learn normal-anomaly boundaries. Our proposed dynamic video synthesis addresses this issue by concatenating semantically similar video segments to reconstruct normal contextual information, with the detailed workflow described as follows.
Algorithm 1 Dynamic Video Synthesis Module
1:Normal video set $N$, Abnormal video set $A$, Synthesis probability $\theta\in[0,1]$, Normal video probability $\alpha\in[0,1]$, Maximum number of segments $\delta_{m}\in\mathbb{N}^{+}$, Number of nearest neighbors $n\in\mathbb{N}^{+}$
2:Synthesized video segment sequence $v$
3:Initialize: video sequence $v$
4:Sample $p_{1},p_{2}\sim U(0,1)$
5:if $p_{1}>\theta$ then
6:  $m\leftarrow 1$
7:else
8:  $m\leftarrow\text{Randint}(1,\delta_{m})$
9:end if
10:if $p_{2}>\alpha$ then
11:  Sample $\tilde{v}$ from AA
12:else
13:  Sample $\tilde{v}$ from NN
14:end if
[…]
16:for $i=1$ to $m$ do
17:  if $i=j$ then
18:   Append $\tilde{v}$ to $v$
19:  else
20:   Sample $v_{n}$ from $\text{NearestNeighbor}(N,\tilde{v},n)$
21:   Append $v_{n}$ to $v$
22:  end if
23:end for
24:return $v$
Figure A: An example of a synthesized video with our proposed dynamic video synthesis.
Dynamic video synthesis comprises a synthesis module and its corresponding pseudo-label loss. Algorithm 1 illustrates the workflow of our synthesis module. The process begins by randomly selecting either a normal or abnormal video as the anchor $\tilde{v}$ (determining whether the synthesis target is normal or abnormal). Subsequently, it randomly determines the number of synthetic segments $m$ and their insertion positions $j$. The remaining positions are then populated by randomly selecting videos from the $N$ nearest neighbors in the normal video set. For nearest neighbor computation, we employ CLIP-ViT-B/16 features extracted from the central frame of each video, using cosine similarity as the distance metric. Unlike the 10-crop approach that pre-generates the augmented data, our method dynamically synthesizes samples during training. To accelerate training efficiency, we pre-compute the $N$ nearest neighbors for each sample in advance.
Since the positions of abnormal segments are known during synthesis, we sequentially construct pseudo-labels $y^{\text{p}}\in\mathbb{R}^{L}$, where $L$ denotes the total length of the synthesized video features. For synthesized normal videos, all elements in $y^{\text{p}}$ are set to 0. For synthesized abnormal videos, $y^{\text{p}}$ takes the value 1 in abnormal intervals and 0 elsewhere. Fig. A shows an example. Using these pseudo-labels, we compute the following loss function:
$-\textstyle\sum_{i}^{L}y^{\text{p}}_{i}\log\sigma(y^{\text{bin}}_{i})/L,$ (16)
where $\sigma$ denotes the Sigmoid function, $\hat{y}$ denotes the video-level ground truth, $L$ denotes the feature length, $\Omega^{a}_{k}$ and $\Omega^{n}_{k}$ are indices of Top-K scores of synthetic abnormal and normal videos, respectively. The first two terms can be viewed as variants of MIL loss that constrain the selection of abnormal instances to regions with pseudo-label value 1, while the last term directly leverages pseudo-labels for supervised learning.
Temporal Encoder
We employ a vanilla Transformer with rotary positional encoding (Su et al., 2024) as our temporal encoder.
Figure B: Detail architecture of fusion module and heads.
Fusion
As illustrated in Fig. B, our vision-text modality fusion adopts a simple co-attention Transformer architecture. Each modality branch contains a cross-attention layer (CA) followed by a feed-forward network (FFN), where the current modality serves as the query in the cross-attention while the other modality provides keys and values.
Heads
As illustrated in Fig. B, the temporal temporal detection head utilizes a 1D convolutional network with replicate padding, which processes both pre-fusion and post-fusion features through separate pathways to generate language-guided and language-agnostic anomaly scores. These scores are subsequently fused via a learnable parameter to produce final predictions. The video-level multi-class classification head computes a similarity matrix between linearly projected representations of both modalities.
B.2 Implementation
All experiments could be conducted on a single NVIDIA RTX4090 GPU. We use PyTorch and PyTorch-Lightning as the code architecture. We set the temperature factors $\tau,\eta$ as 0.02 and employ a base hidden size of 512. The temporal encoder and the fusion module both have 2 layers, and the detection head uses a single layer with a kernel size of 9. For dynamic video synthesis, we set $\theta=0.7,\alpha=0.5,\delta_{m}=5,n=200$. We use AdamW as the optimizer with a batch size of 64 and a learning rate of 0.00005. The model is trained for 40 epochs. We do not use tricks like 5-crop, 10-crop, or score smooth, etc.. We take a sample every 8 frames, except for the DoTA dataset, which provides 10fps extracted frames.
Figure C: The data curation pipeline for our dataset. It includes three phases: data source, cleansing, and annotation.
Table A: Details about the data sources used by PreVAD.
Source Notes
video-text datasets VIDAL-10M (Zhu et al., 2024a)
The VIDAL-10M dataset comprises 3 million pairs of video-language data crawled from
YouTube Shorts and Freesound. It includes abnormal videos of accidents and normal videos
of everyday events taken by people with their smart phones.
MSR-VTT (Xu et al., 2016)
The MSR-VTT dataset is a video captioning dataset that comprises 10K videos
with 20 captions per video. The videos are crawled from YouTube and are annotated
by crowdsourcing. It contains many domains, including human activities, tutorials, games,
etc. We mainly use them as normal data.
VATEX (Wang et al., 2019)
The VATEX dataset is also a video captioning dataset with 35K videos,
each having 10 manually labeled captions. Its videos are from the Kinetics dataset,
and thus are mostly diverse human activities. We mainly use them as normal data.
VALOR (Liu et al., 2025)
The VALOR dataset is a video captioning dataset that focuses on audio-based description.
The authors release a subset of 32K videos, each with one manually annotated caption.
The videos are from AudioSet, which covers human activities and natural scenes.
We mainly use them as normal data.
web videos collections
There are dedicated channels on YouTube and Bilibili that compile a variety of
accident videos for the purposes of entertainment or education (e.g., FailsArmy, ASP).
We clip segments from them as abnormal data.
long videos
Some channels post hour-long videos, the content of which includes
first-person-perspective driving, first-person-perspective traveling, and factory
production records, etc. We randomly clip segments from such videos as normal data.
(e.g., Driving)
RWF-2000 (Cheng et al., 2020)
The RWF-2000 dataset is a violence recognition dataset that includes violent behaviors
from a surveillance perspective, and we incorporate it into our dataset.
streaming YouTube streams
On YouTube, there are numerous 24-hour surveillance live streams aimed at promoting tourism.
We searched for such content using the keyword webcam and recorded segments
at different times as normal data from a surveillance perspective. (e.g., Manchester UK Webcams)
traffic cameras
With the development of intelligent transportation, many countries have deployed
traffic cameras on highways and made their real-time footage publicly available
on the Internet to help the public better plan their travel. We also recorded these
as normal data. (e.g.,California Department of Transportation,
Shanghai Municipal Transportation Commission)
Appendix C Details of PreVAD
C.1 Dataset Statement
The dataset introduced in this paper is an original collection compiled by the authors for the purpose of advancing research in video anomaly detection. All data within the dataset are obtained from public sources, and there are no privacy and ethical issues involved. The dataset is intended for academic and research use only. Any commercial use or unauthorized distribution is strictly prohibited without the prior written consent of the authors. We release the annotation of the dataset and scripts for download.
C.2 Pipeline
Fig. C illustrates the proposed data curation pipeline, leveraging publicly accessible sources for scalability and reproducibility, integrating multimodal foundation models (LLMs, MLLMs, VLMs) for intelligent cleaning, and employing a hybrid human-AI annotation framework to ensure quality while reducing costs.
Source
Tab. A lists the data source details for PreVAD. An important issue faced by video anomaly detection for many years is data scarcity, and we exploit multiple data sources. To the best of our knowledge, we are the first to enumerate and analyze the data sources in detail. We mainly collected the data in 2024. All data sources are publicly available. Using these datasets, future works may construct a larger scale dataset.
Cleansing
During the text-based video retrieval, we employ all-MiniLM-L6-v2 and CLIP-Large Text Encoder to extract text embeddings of a set of abnormal queries (e.g., a fire broke out in the building) and captions in existing video-text datasets as keys. Then, we use FAISS (Douze et al., 2024) to retrieve abnormal captions and obtain their corresponding videos. We preliminarily retrieved over 55K videos in this phase, among which only nearly 75% of videos are available when downloading. During the irrelevant clips removal, we employ PySceneDetect111github.com/Breakthrough/PySceneDetect to split segments of the video and then automatically filter them with predefined rules of duration, RGB histogram, etc.. During the MLLM-based video captioning phase, we prompt Qwen2-VL-72B-Instruct with:
You are a helpful video describing assistant. You are good at English communication. Please describe the given video in detail. Just give the description without any other output.
to obtain captions of existing videos for later filtering. During the VLM-based video-text matching, we utilize InternVideo2 (Wang et al., 2024b) and VALOR (Liu et al., 2025) to calculate the matching scores of video-text pairs to remove irrelevant pairs with a threshold. During the LLM-based abnormality judging, we prompt Qwen/Qwen2-72B-Instruct with:
Below is the JSON format metadata of the title and content information of a video. I need you to determine if there are any anomalies in the video. Anomalies include:
- any type of accident
Note, you need to output two lines, the first line is a brief one-sentence analysis, and the second line is a judgment of ‘abnormal’ or ‘normal’. If you cannot provide analysis due to ethical guidelines, please just output ‘abnormal’.
to filter videos with wrong abnormality labels. After cleansing, we could obtain videos and corresponding video-level binary labels.
Annotation
During the human annotation phase, we asked 7 annotators with expert knowledge for video anomaly detection to label each abnormal video with one video-level category among our taxonomy. They are also responsible for removing videos of low quality (e.g., anomalies far from the real world, excessive video effects or overlays of text, excessively short videos, and videos containing clips of multiple different scenes). For the validation set, we let two annotators independently annotate the segments of the selected abnormal videos and take the average as the final result. If the annotation difference between the two people is too large, a team leader will recheck. During the MLLM-based prompted abnormal video describing phase, we prompt Qwen2-VL-72B-Instruct with:
You are a helpful video describing assistant. You are good at English communication. Please describe the ‘{VIDEO LABEL}‘ event in the given video in detail in one sentence. The description should focus on the incident. Just describe without any other output.
to obtain precise anomaly description. With this constrained prompt, the descriptions would be more related to the anomaly, therefore, are suitable for being anomaly definitions.
(a) The multi-level taxonomy of PreVAD.
(b) The treemap of PreVAD’s category distribution.
(c) Distribution of coarse-grained categories.
(d) Detailed distribution of video sources.
(e) Distribution of the proportion of anomaly in the validation set.
(f) Distribution of the duration.
Figure D: More statistics of PreVAD.
C.3 More Statistics
Fig. 4(a) provides the taxonomy of PreVAD, with Fig. 4(b) and Fig. 4(c) show the distribution of categories. Fig. 4(d) provides a detailed source distribution visualization. Fig. 4(e) shows the proportion of anomaly in annotated abnormal videos and Fig. 4(f) presents the distribution of video duration in seconds.
C.4 Examples
We show examples of videos for each category and annotated examples in Figs. F,G. Our data contains a variety of high-quality anomaly videos, and the human-AI annotation pipeline can obtain accurate anomaly descriptions at a low cost.
Appendix D Details of Evaluation
D.1 Test Sets
Tab. B summarizes the number of test videos, anomaly categories, and other relevant details for the seven datasets used as test sets. These datasets cover a diverse range of anomalies across domains such as crime (UCF-Crime, XD-Violence, LAD, MSAD), pedestrian (UBNormal), and traffic (DoTA, LAD), where each test set represents a distinct scenario with its own anomaly definition. There exists many categories that are not included in PreVAD, where some are unseen categories (e.g., Arrest), while others appear in PreVAD’s normal videos (e.g., Pedestrian on Road). This poses a challenge to the open-world generalization ability of the model. Notably, the DoTA dataset only contains abnormal videos, therefore, the predictions on this dataset are normalized for evaluation.
Table B: Overview of test datasets. Abnormal categories not included in PreVAD are underlined: some are unseen categories (e.g., Arrest), while others appear in PreVAD’s normal videos (e.g., Pedestrian on Road).
Datasets # testing videos Categories Notes
UCF-Crime 290
Abuse, Arrest, Arson, Assault, Accident, Burglary,
Explosion, Fighting, Robbery, Shooting, Stealing, Shoplifting, Vandalism
crime scene from surveillance camera
XD-Violence 800 Abuse, Car accident, Explosion, Fighting, Riot, Shooting
videos contain shot cut, and part of
the data comes from movies
MSAD 240
Assault, Fighting, People Falling, Robbery, Shooting, Traffic Accident,
Vandalism, Explosion, Fire, Object Falling, Water Incident
the accident in the perspective
UBNormal 211
running, having a seizure, laying down, shuffling, walking drunk,
people and car accident, car crash, jumping, fire, smoke,
jaywalking, driving outside lane
videos of 3D animation
DoTA 1402 (9 kinds of fine-grained traffic anomalies)
only contains abnormal videos in
both first-person and third-person view
TAD 100
Accidents, Illegal Turns, Illegal Occupations,
Retrograde Motion, Pedestrian on Road, Road Spills, Else
abnormal videos of traffic in both
first-person and third-person view
LAD 560
Crash, Crowd, Destroy, Drop, Falling, Fighting, Fire,
Fall Into Water, Hurt, Loitering, Panic, Thiefing, Trampled, Violence
accidents from surveillance and
first-person perspectives
D.2 Protocols
We set up two protocols to evaluate the model, where Protocol 1 comprehensively evaluates the open-world generalization capability of models, and Protocol 2 specifically evaluates robustness to concept drift.
Protocol 1
Models are trained on pretraining datasets (e.g., PreVAD) and evaluated in a zero-shot manner across multiple test sets, each representing a distinct scenario with unique definitions of anomalies, video distributions, and anomaly categories. For instance, UCF-Crime focuses on crime-related scenarios and contains normal videos showing pedestrians crossing roads—a behavior considered abnormal in TAD, which focuses on traffic scenarios. Additionally, while UCF-Crime and MSAD primarily consist of surveillance footage, XD-Violence and DoTA include movie clips and dashcam videos, respectively. According to Tab. B, this protocol simulates the actual open-world situation, i.e., training on limited data and testing in different scenarios.
Protocol 2
Models are trained on pretraining datasets (e.g., PreVAD) and tested in a zero-shot setting under varying anomaly definitions. Each definition corresponds to a randomly selected subset of anomaly categories from the test set. Videos belonging to the chosen subset are treated as abnormal, while those outside the subset are considered normal. To ensure reliability, we randomly sample five subsets per test set and report the average performance (denoted as drift@5). The selected subsets are detailed in Tab. C. For example, when subset #1 is chosen, videos labeled as fighting, shooting, and riot are regarded as abnormal, whereas videos labeled as normal, abuse, car accident, and explosion are considered normal. This protocol is specifically designed to assess the ability to handle concept drift issue, which is the main focus of this paper.
Table C: Drift@5 settings for evaluation protocol 2.
Subset XD-Violence MSAD
1 Fighting, Shooting, Riot Assault, Fighting, Robbery, Shooting, Vandalism, Explosion
2 Abuse, Car accident, Explosion People_falling, Traffic_accident, Fire, Object_falling, Water_incident, Explosion
3 Fighting, Abuse, Explosion Assault, People_falling, Traffic_accident, Vandalism, Fire, Object_falling
4 Shooting, Car accident, Riot Fighting, Robbery, Shooting, Traffic_accident, Water_incident, Fire
5 Riot, Abuse, Car accident People_falling, Robbery, Object_falling, Water_incident, Explosion, Assault
Table D: Ablations of hyperparameters in dynamic video synthesis. Following symbols in Algorithm 1, $\delta_{m}$ denotes the max number of segments, $\theta$ denotes the synthesis probability, and KNN refers to retrieval of K nearest neighbors.
$\delta_{m}$ $\theta$ KNN PreVAD UCF-Crime XD-Violence
5 0.7 69.98 81.12 74.25
1 - 65.73 79.18 71.41
3 0.7 68.24 80.34 67.96
7 0.7 66.99 80.60 78.09
9 0.7 67.13 79.28 72.90
5 0.7 67.85 79.98 68.95
5 1 67.86 78.81 71.63
5 0.3 66.53 80.50 68.74
Appendix E Details of Reproduced Methods
Since this is a pioneering project, most of our comparisons are based on the results we reproduced with open-source codes and weights. The compared methods can be divided into two groups: non-LLM-based and LLM-based approaches.
Non-LLM-based methods
We compare PEL (Pu et al., 2024) and VadCLIP (Wu et al., 2024c), which achieve state-of-the-art performance in video anomaly detection, along with ActionCLIP (Wang et al., 2021) and ViFi-CLIP (Rasheed et al., 2023), SOTA methods in open-vocabulary action recognition. Additionally, a simple CLIP baseline is also included. All methods utilize the same ViT-B/16 features, except for PEL, which uses I3D features. For VadCLIP, we employ its A-branch and apply the same post-processing method as ours to obtain classification results. For the CLIP baseline, we uniformly sampled 8 frames from normal videos and 8 frames from abnormal segments of abnormal videos. Similarity are computed between the visual features and textual prompts constructed from category names. The probabilities from the 8 frames are averaged to produce the final prediction. ActionCLIP and ViFi-CLIP are handled similarly to the CLIP baseline, except that they directly process 8-frame video clips and output final results. Although this comparison setup inherently favors these models (as they do not require temporal localization), experimental results demonstrate that our method still achieves superior performance.
LLM-based methods
We compare general-purpose grounding-capable models including Qwen2-VL (Wang et al., 2024a) and Qwen2.5-VL (Bai et al., 2025), the fine-tuned MLLM-based video anomaly detection method HolmesVAU (Zhang et al., 2025), and the multi-VLM ensemble method LAVAD (Zanella et al., 2024). These methods are capable of processing textual prompts, which we use to inject anomaly definition information. Since Qwen2-VL and Qwen2.5-VL cannot output frame-level scores and only predict boundaries, we set the anomaly score of each frame to 0 when no segment is predicted and to 1 for frames within predicted boundaries. HolmesVAU could not be included in the zero-shot comparison on XD-Violence since its training set contains videos from XD-Violence. Although a direct comparison of computational efficiency is challenging due to differences in hardware and implementation optimizations, our proposed method are significantly faster (minutes vs. hours during evaluation).
Table E: Ablation on prompting method during inference.
Prompt method UCF-Crime AUC
class name 80.44
manual prompt (default) 81.12
video-specific description 83.03
Figure E: Ablations of data usage during training.
Table F: Ablation on each component (full version). guided refers to guiding the detection with language. Det. Avg. refers to the average zero-shot detection performance on seven datasets. Cls. Avg. refers to the average zero-shot classification performance on UCF-Crime and XD-Violence.
$\mathcal{L}_{\text{dvs}}$ $\mathcal{L}_{\text{neg}}$ guided PreVAD average metrics Detection Classification
Det. Avg. Cls. Avg. UCF XD MSAD UBN DoTA TAD LAD UCF ACC UCF F1 XD ACC XD F1
69.98 76.42 52.57 81.12 74.25 90.41 58.07 62.60 89.56 78.91 51.72 16.64 78.13 63.80
65.73 73.51 51.73 79.18 70.24 84.14 55.79 60.02 85.95 79.23 44.83 16.25 79.00 66.82
68.92 73.96 51.85 77.66 70.66 90.02 57.56 59.88 83.26 78.65 48.97 17.76 75.88 64.78
67.35 71.31 48.81 77.77 62.49 89.50 55.80 57.50 80.92 75.19 43.82 17.60 73.75 60.06
69.87 73.84 46.23 76.64 71.39 88.60 56.88 61.40 83.96 78.02 46.21 13.28 66.87 58.54
Appendix F More Experimental Results
F.1 Ablations
In Tab. E, we experiment with different prompting methods: 1) class name uses only category labels as anomaly definitions (e.g., explosion, fighting), 2) manual prompt employs human-designed description of the category (e.g., Explosion, often resulting in fire, smoke, and scattered debris), which serves as our default approach, and 3) video-specific description utilizes the description of specific abnormal events as definitions (e.g., two people catch fire in the explosion near the garage). Experimental results demonstrated that better prompts could enhance performance. Although video-specific descriptions are unavailable in standard test datasets, it reveals our method’s potential for practical applications like locating relevant surveillance clips when specific event details are known.
We provide the complete version of Tab. 5 in Tab. F, which contains more metrics of ablations. We also report hyperparameter search results of dynamic video synthesis in Tab. D. And Tab. E illustrates the importance of large-scale datasets.
F.2 Qualitative Results
As in Fig. H, we present more zero-shot comparisons on different datasets. The compared three methods all performed visual-language alignment. VadCLIP and LaGoVAD, which utilize pre-aligned feature extractors, demonstrate a significant advantage in cross-dataset testing, further proving the importance of semantic alignment for generalization. Additionally, our proposed LaGoVAD achieves predictions with fewer false alarms, thanks to the contrastive learning with in-sample negative mining we conducted. Finally, LaGoVAD was also able to make accurate predictions for the novel Arrest category.
Fig. I shows a prediction example of LaGoVAD on the XD-Violence dataset. The primary anomaly in this video is a riot scene, accompanied by secondary anomalies including fighting, explosion, and shooting. We visualize the outputs from both the detection head and the classification head. Our model achieves accurate detection in normal intervals (a-c) as well as in riot scenarios (b). Furthermore, the model effectively captures mixed fighting (e) and shooting events (f-g) through its classification head, demonstrating a sensitive response to co-occurring anomalies. Notably, even for transient explosion events (d) with extremely short durations, our model could produce sharp response spikes.
Figure F: Examples of each abnormal category in PreVAD.
Figure G: Examples of annotations in PreVAD.
Figure H: Qualitative comparisons with other models. The two models used for comparison are trained on the UCF dataset (a)(b)(c) or the XD-Violence dataset (d). All the results are zero-shot results.
Figure I: Visualization of results of the detection head and the classification head of the proposed LaGoVAD. The top-left plot illustrates the detection head output, while the remaining plots correspond to the classification head outputs, activated via the Sigmoid function. Frames (a)-(g) denote key frames. LaGoVAD is able to detect fine-grained anomalies in a video that contains multiple different anomalies.